# Non-Homophilic Graph Pre-Training and Prompt Learning

### v4

@@ 4.2. Non-homophilic Graph Pre-training
Consider a homophily task $T$. According to Theorem 4, for non-homophilic graphs with lower homophily ratios, on average there are fewer homophily samples and more non-homophily samples for $T$. Consequently, based on Theorem 3, adding a non-homophily sample always results in a larger loss than adding a homophily sample. Therefore, for non-homophilic graphs, especially those with low homophily ratio, non-homophily tasks are a better choice compared to homophily tasks when optimizing the training loss.
We revisit mainstream graph pre-training methods and categorize them into two categories: homophily methods that employ homophily tasks, and non-homophily methods that do not. Specifically, GMI (Peng et al., 2020), GraphPrompt (Liu et al., 2023a), MultiGPrompt (Yu et al., 2024d), HGPrompt (Yu et al., 2024b) and GraphPrompt+ (Yu et al., 2024c) are all homophily methods, since their pre-training tasks utilizes a form of link prediction, where $\mathcal{A}_{u}$ is a set of nodes linked to $u$, and $\mathcal{B}_{u}$ is a set of nodes not linked from $u$. In contrast, DGI (Velickovic et al., 2019), GraphCL (You et al., 2020), and GraphACL (Xiao et al., 2023) are non-homophily methods, since $\mathcal{A}_{u}$ and $\mathcal{B}_{u}$ in their pre-training tasks are not related to the connectivity with $u$. Further details of these methods are shown in Appendix. B. In our experiment, we use the non-homophily method GraphCL as the pretext task to obtain our main results for non-homophily graphs, as it is a classic pre-training method with competitive performance. We also experiment with link prediction (Liu et al., 2023a) and GraphACL for further evaluation, as shown in Table 4.
@@ 5. Non-homophilic Prompt Learning
In this section, we propose ProNoG, our prompt learning framework for non-homophilic graphs. We first introduce the overall framework, and then develop the prompt generation and tuning process. Finally, we present the overall algorithm and analyze its complexity.
@@ 5.2. Prompt Generation and Tuning
(7) $\mathbf{s}_{v}=\frac{1}{|S_{v}|}\sum_{u\in S_{v}}\mathbf{h}_{u}\cdot\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{v}),$
where $|S_{v}|$ denotes the number of nodes in $S_{v}$. In our experiment, we set $\delta=2$ to balance between efficiency and capturing more unique non-homophilic patterns in the neighborhood of $v$.
@@
where $\mathbf{\tilde{h}}_{t,x_{i}}$ denotes the output embedding of node $v$/graph $G$ for task $t$. Specifically, for node classification $\mathbf{\tilde{h}}_{t,{v}}$ is the output embedding in Eq. 9; for graph classification, $\mathbf{\tilde{h}}_{t,{G}}=\sum_{u\in V}\mathbf{\tilde{h}}_{t,{u}}$ involving an additional graph readout. The prototype embedding for class $c$, $\mathbf{\bar{h}}_{t,c}$, is the average of the output embedding of all nodes/graphs belonging to class $c$.
During prompt tuning, we update only the lightweight parameters of the condition-net ($\phi_{t}$), while freezing the pre-trained GNN weights. Thus, our prompt tuning is parameter-efficient and amenable to few-shot settings, where $\mathcal{D}_{t}$ contains only a small number of training examples for task $t$.
@@ 5.3. Algorithm and Complexity Analysis
Algorithm. We detail the main steps for conditional prompt generation and tuning in Algorithm 1, Appendix A.
Complexity analysis. For a downstream graph $G$, the computational process of ProNoG involves two main parts: encoding nodes via a pre-trained GNN, and conditional prompt learning. The first part’s complexity is determined by the GNN’s architecture, akin to other methods employing a pre-trained GNN. In a standard GNN, each node aggregates features from up to $n$ neighbors per layer. Assuming the aggregation involves at most $D$ neighbors, the complexity of calculating node embeddings over $L$ layers is $O(D^{L}\cdot|V|)$, where $|V|$ denotes the number of nodes. The second part, conditional prompt learning, has two stages: prompt generation and prompt tuning. In the prompt generation stage, each subgraph embedding is fed into the condtion-net. In our experment, we use a 2 layer MLP as condition-net, resulting in a complexity of $O(2\cdot|V|)$. During prompt tuning, each node in $G$ is adjusted using a prompt vector, with a complexity of $O(|V|)$. Therefore, the total complexity for conditional prompt learning is $O(3\cdot|V|)$.
In conclusion, the overall complexity of ProNoG is $O((D^{L}+3)\cdot|V|)$. The first part dominates the overall complexity, as $O(n^{L}\cdot|V|)\gg O(3\cdot|V|)$, where we set $L=2$ for experiments on low-homophily graphs. Thus, the additional computational cost introduced by the conditional prompt tuning step is minimal.
@@ 6.1. Experimental Setup
Datasets. We conduct experiments on ten benchmark datasets. Wisconsin (Pei et al., 2020), Cornell (Pei et al., 2020), Chameleon (Rozemberczki et al., 2021), and Squirrel (Rozemberczki et al., 2021) are all webpage graphs. Each dataset features a single graph where nodes correspond to web pages and edges represent hyperlinks connecting these pages. Cora (McCallum et al., 2000) and Citeseer (Sen et al., 2008) are citation networks. These datasets consist of a single graph each, with nodes signifying scientific papers and edges indicating citation relationships. PROTEINS (Borgwardt et al., 2005) consists of a series of protein graphs. Nodes in these graphs denote secondary structures, while edges depict neighboring relationships either within the amino acid sequence or in three-dimensional space. ENZYMES (Wang et al., 2022), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are collections of molecular graphs. These datasets describe enzyme structures from the BRENDA enzyme database, ligands related to benzodiazepine receptors, and cyclooxygenase-2 inhibitors, respectively. We summarize these datasets in Table 6, and present further details in Appendix C.
Baselines. We evaluate ProNoG against a series of state-of-the-art methods, categorized into three primary groups:
(1) End-to-end graph neural networks: GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), H2GCN (Zhu et al., 2020b), and FAGCN (Bo et al., 2021) are trained in a supervised manner directly using downstream labels. Specifically, GCN and GAT are originally designed for homophilic graphs, H2GCN for heterophilic graphs, and FAGCN for non-homophilic graphs.
(2) Graph pre-training models: DGI (Velickovic et al., 2019), GraphCL (You et al., 2020), DSSL (Xiao et al., 2022), GraphACL (Xiao et al., 2023) follow the “pre-train, fine-tune” paradigm.
(3) Graph prompt learning models: GPPT (Sun et al., 2022), GraphPrompt (Liu et al., 2023a), and GraphPrompt+ (Yu et al., 2024c) use self-supervised pre-training tasks with a single type of prompt for downstream adaptation. Note that GPPT is specifically designed for node classification and cannot be directly used for graph classification. Therefore, in our experiments, we use GPPT exclusively for node classification tasks.
@@
We provide further details on these baselines in Appendix D. It’s worth noting that some graph few-shot learning methods, such as Meta-GNN (Zhou et al., 2019), AMM-GNN (Wang et al., 2020), RALE (Liu et al., 2021a), VNT (Tan et al., 2023), and ProG (Sun et al., 2023a), are based on the meta-learning paradigm (Finn et al., 2017), which requires an additional set of labeled base classes in addition to the few-shot classes. Consequently, these methods are not directly comparable to our framework.
@@
Parameter settings. For all baselines, we use the original authors’ code and follow their recommended settings, while further tuning their hyperparameters to ensure optimal performance. Detailed descriptions of the implementations and settings for both the baselines and our ProNoG are provided in Appendix E.
@@
Setup of downstream tasks. We conduct two types of downstream task: node classification, and graph classification. These tasks are set up as $k$-shot classification problems, meaning that for each class, $k$ instances (nodes or graphs) are randomly selected for supervision. Given that all low-homophily datasets, i.e., Wisconsin, Squirrel, Chameleon and Cornell only comprise a single graph and cannot be directly used for graph classification. Thus, following previous research (Lu et al., 2021; Yu et al., 2024b), we generate multiple graphs by constructing ego-networks centered on the labeled nodes in each dataset. We then perform graph classification on these ego-networks, each labeled according to its central node. For datasets with high homophily ratios, PROTEINS, ENZYMES, BZR and COX2 have original graph labels, so we directly conduct graph classification on these graphs. Since the $k$-shot tasks are balanced classification problems, we use accuracy to evaluate performance, in line with prior studies (Liu et al., 2023a; Wang et al., 2020; Liu et al., 2021a; Yu et al., 2024c). We pre-train the graph encoder once for each dataset and then use the same pre-trained model for all downstream tasks. We generate 100 $k$-shot tasks for both node classification and graph classification by repeating the sampling process 100 times. Each task is executed with five different random seeds, leading to a total of 500 results per task type. We report the mean and standard deviation of these 500 outcomes.
@@
Table 4. Comparison between homophily and non-homophily pre-training.
@@ 6.2. Few-shot Performance Evaluation
One-shot performance. We present the results of one-shot node and graph classification tasks on non-homophilic graphs in Tables 1 and 2, respectively. We make the following observations: (1) ProNoG surpasses all baseline methods across all settings, outperforming the best competitor by up to 21.49% on node classification and 6.50% on graph classification. These results demonstrate its effectiveness in learning prior knowledge from non-homophilic graphs and capturing nodes’ specific patterns. (2) Other graph prompt learning methods, i.e., GPPT, GraphPrompt, and GraphPrompt+, significantly lag behind ProNoG. Their suboptimal performance can be attributed to their inability to account for a variety of node-specific patterns. These results underscore the importance of our conditional prompting in characterizing node embeddings to capture nodes’ specific patterns. (3) GPPT is at best comparable to, and often performs worse than other baselines because it is not specifically designed for few-shot learning.
Few-shot performance. To assess the performance of ProNoG with different amounts of labeled data, we vary the number of shots in the downstream tasks and present the results in Fig. 3 and Appendix F. Note that given the limited number of nodes in Wisconsin and Cornell, we only conduct tasks up to $3$-shot. We observe that: (1) ProNoG significantly outperforms all baselines in low-shot scenarios with very limited labeled data (e.g., $k\leq 5$), showcasing the effectiveness of our approach in these situations. (2) As the number of shots increases, all methods generally show improved performance as expected. However, ProNoG remains competitive and often surpasses the other methods, demonstrating the robustness of ProNoG.
@@
Figure 4. Results on different node patterns.
@@ 6.4. Analysis on Pre-Training Methods
To further evaluate homophily and non-homophily tasks, using ProNoG for downstream adaptation, we employ homophily tasks link prediction used in GraphPrompt (Liu et al., 2023a), and non-homophily tasks GraphCL (You et al., 2020) and DSSL (Xiao et al., 2022), respectively. Note that link predicition in GPPT (Sun et al., 2022) is in a generative format, thus falling beyond the scope of homophily task, but it’s also affected by non-homophily in graphs. We compare these pretext tasks and show the results in Table 4. We observe that for graphs with a low homophily ratio, the non-homophily task significantly outperforms the homophily tasks. Conversely, for graphs with a high homophily ratio, the results of these two methods are mixed, with each having their own strengths and weaknesses.
6.5. Analysis on Varying Node Patterns
To evaluate the ability of ProNoG in capturing node-specific patterns, we calculate the accuracy on different node groups with varying homophily ratios, i.e., $[0.0,0.2)$, $[0.2,0.4)$, $[0.4,0.6)$, $[0.6,0.8)$, $[0.8,1.0]$. We compare ProNoG with several competitive baselines and present the results in Fig. 4. We observe that ProNoG consistently outperforms or is at least competitive with all baselines across all node patterns, regardless of their homophily ratio. These results further demonstrate the effectiveness of ProNoG in capturing node-specific patterns and highlight the advantages of our proposed conditional prompting.
@@ 7. Conclusions
In this paper, we explored pre-training and prompt learning on non-homophilic graphs. The objectives are twofold: learning comprehensive knowledge irrespective of the varying non-homophily characteristics of graphs, and adapting the nodes with diverse distributions of non-homophily patterns to downstream applications in a fine-grained, node-wise manner. We first revisit graph pre-training on non-homophilic graphs, providing theoretical insights into the choice of pre-training tasks. Then, for downstream adaptation, we proposed condition-net to generate a series of prompts conditioned on various non-homophilic patterns across nodes. Finally, we conducted extensive experiments on ten public datasets, demonstrating that ProNoG significantly outperforms diverse state-of-the-art baselines.

### v6

@@ 4.2. Non-homophilic Graph Pre-training
Consider a homophily task $T$. Following Theorem 4, non-homophilic graphs with lower homophily ratios are expected to have fewer homophily samples and more non-homophily samples for $T$. Furthermore, based on Theorem 3, adding a non-homophily sample results in a larger loss than adding a homophily sample. Consequently, for non-homophilic graphs, especially those with low homophily ratio, homophily tasks are not optimal for working with standard contrastive training losses, whereas non-homophily tasks may offer a better alternative.
We revisit mainstream graph pre-training methods and categorize them into two categories: homophily methods that employ homophily tasks, and non-homophily methods that do not. Specifically, GPPT (Sun et al., 2022), GraphPrompt (Liu et al., 2023a) and HGPrompt (Yu et al., 2024b) are all homophily methods, since their pre-training tasks utilizes a form of link prediction, where $\mathcal{A}_{u}$ is a set of nodes linked to $u$, and $\mathcal{B}_{u}$ is a set of nodes not linked from $u$. In contrast, DGI (Velickovic et al., 2019), GraphCL (You et al., 2020), and GraphACL (Xiao et al., 2023) are non-homophily methods, since $\mathcal{A}_{u}$ or $\mathcal{B}_{u}$ in their pre-training tasks is not related to the connectivity with $u$. Further details of these methods are shown in Appendix B.
In our experiments, we find that GraphCL (You et al., 2020), a non-homophily pre-training method, performs well across most graphs, including non-homophilic ones. We also experiment with link prediction (Liu et al., 2023a; Sun et al., 2022) and GraphACL (Xiao et al., 2023) to study their effects on different graphs.
@@ 5. Non-homophilic Prompt Learning
In this section, we propose ProNoG, a prompt learning framework for non-homophilic graphs. We first introduce the overall framework, and then develop the prompt generation and tuning process. Finally, we analyze the complexity of the proposed algorithm.
@@ 5.2. Prompt Generation and Tuning
(7) $\mathbf{s}_{v}=\frac{1}{|V(S_{v})|}\sum_{u\in V(S_{v})}\mathbf{h}_{u}\cdot\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{v}),$
where $V(S_{v})$ denotes the set of nodes in $S_{v}$. In our experiment, we set $\delta=2$ to balance between efficiency and capturing more unique non-homophilic patterns in the neighborhood of $v$.
@@
where $\mathbf{\tilde{h}}_{t,x_{i}}$ denotes the output embedding of node $v$/graph $G$ for task $t$. Specifically, for node classification $\mathbf{\tilde{h}}_{t,{v}}$ is the output embedding in Eq. (9); for graph classification, $\mathbf{\tilde{h}}_{t,{G}}=\sum_{u\in V}\mathbf{\tilde{h}}_{t,{u}}$, involving an additional graph readout. The prototype embedding for class $c$, $\mathbf{\bar{h}}_{t,c}$, is the average of the embedding of all labeled nodes/graphs belonging to class $c$.
During prompt tuning, we update only the lightweight parameters of the condition-net ($\phi_{t}$), while freezing the pre-trained GNN weights. Thus, our approach is parameter-efficient and amenable to few-shot settings, where $\mathcal{D}_{t}$ contains only a small number of training examples for task $t$.
@@ 5.3. Algorithm and Complexity Analysis
Algorithm. We detail the main steps for the conditional prompt generation and tuning in Algorithm 1, Appendix A.
Complexity analysis. For a downstream graph $G$, the computational process of ProNoG involves two main parts: encoding nodes via a pre-trained GNN, and conditional prompt learning. The first part’s complexity is determined by the GNN architecture, akin to other methods employing a pre-trained GNN. In a standard GNN, each node aggregates features from up to $D$ neighbors per layer. Thus, the complexity of calculating node embeddings over $L$ layers is $O(D^{L}\cdot|V|)$, where $|V|$ denotes the number of nodes. The second part, conditional prompt learning, has two stages: prompt generation and prompt tuning. In prompt generation, each subgraph embedding is fed into the condition-net. The subgraph embedding of each node involves a readout from the $\delta$-hop neighborhood, resulting in a complexity of $O(D^{\delta}\cdot|V|)$ with at most $D$ neighbors per hop. During prompt tuning, each node in $G$ is adjusted using a prompt vector, with a complexity of $O(|V|)$. Therefore, the total complexity for conditional prompt learning is $O(D^{\delta}\cdot|V|)$.
In conclusion, the overall complexity of ProNoG is $O((D^{L}+D^{\delta})\cdot|V|)$. As both $L$ and $\delta$ are small constants, the two parts have comparable complexity. That is, the proposed conditional prompt learning does not increase the order of complexity relative to the pre-trained GNN encoder, if $\delta$ is chosen to be no larger than $L$.
@@ 6.1. Experimental Setup
Datasets. We conduct experiments on ten benchmark datasets. Wisconsin (Pei et al., 2020), Cornell (Pei et al., 2020), Chameleon (Rozemberczki et al., 2021), and Squirrel (Rozemberczki et al., 2021) are all web graphs. Each dataset features a single graph, where nodes correspond to web pages and edges represent hyperlinks connecting these pages. Cora (McCallum et al., 2000) and Citeseer (Sen et al., 2008) are citation networks. They consist of a single graph each, with nodes representing academic papers and edges indicating citation relationships. PROTEINS (Borgwardt et al., 2005) consists of a series of protein graphs. Nodes in these graphs denote secondary structures, while edges depict neighboring relationships either within the amino acid sequence or in three-dimensional space. ENZYMES (Wang et al., 2022), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are collections of molecular graphs. These datasets describe enzyme structures from the BRENDA enzyme database, ligands related to benzodiazepine receptors, and cyclooxygenase-2 inhibitors, respectively. We summarize these datasets in Table 6, Appendix C.
Baselines. We evaluate ProNoG against a series of state-of-the-art methods from the following three categories. (1) End-to-end GNNs: GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), H2GCN (Zhu et al., 2020b), and FAGCN (Bo et al., 2021) are trained in a supervised manner directly using downstream labels. Specifically, GCN and GAT are designed for homophilic graphs, whereas H2GCN is developed for heterophilic graphs, and FAGCN for non-homophilic graphs. (2) Graph pre-training models: DGI (Velickovic et al., 2019), GraphCL (You et al., 2020), DSSL (Xiao et al., 2022), and GraphACL (Xiao et al., 2023) follow the “pre-train, fine-tune” paradigm. (3) Graph prompt learning models: GPPT (Sun et al., 2022), GraphPrompt (Liu et al., 2023a), and GraphPrompt+ (Yu et al., 2024c) employ self-supervised pre-training tasks, and use the same prompts for all nodes in downstream adaptation. Note that GPPT is specifically designed for node classification and cannot be directly used for graph classification. Therefore, we evaluate GPPT on node classification tasks only.
@@
Note that some graph few-shot learning methods, such as Meta-GNN (Zhou et al., 2019), AMM-GNN (Wang et al., 2020), RALE (Liu et al., 2021a), VNT (Tan et al., 2023), and ProG (Sun et al., 2023a), are based on the meta-learning paradigm (Finn et al., 2017). They require a set of labeled base classes in addition to the few-shot classes, and thus are not compared here.
@@
Parameter settings. For all baselines, we use the original authors’ code and reference their recommended settings, while further tuning their hyperparameters to ensure optimal performance. Detailed descriptions of the implementations and settings for both the baselines and our ProNoG are provided in Appendix C.
@@
Downstream tasks and evaluation. We conduct two types of downstream task: node classification, and graph classification. These tasks are set up as $k$-shot classification problems, meaning that for each class, $k$ instances (nodes or graphs) are randomly selected for supervision. The low-homophily datasets, i.e., Wisconsin, Squirrel, Chameleon and Cornell, only comprise a single graph and cannot be directly used for graph classification. Thus, following previous research (Lu et al., 2021; Yu et al., 2024b), we generate multiple graphs by constructing ego-networks centered on the labeled nodes in each dataset. We then perform graph classification on these ego-networks, each labeled according to its ego node. Among datasets with high homophily ratios, PROTEINS, ENZYMES, BZR and COX2 have ground-truth graph labels, which we employ directly for graph classification.
Since the $k$-shot tasks are balanced classification problems, we use accuracy to evaluate the performance, in line with prior studies (Liu et al., 2023a; Wang et al., 2020; Liu et al., 2021a; Yu et al., 2024c). We pre-train the graph encoder once for each dataset and then use the same pre-trained model for all downstream tasks. We generate 100 $k$-shot tasks for both node classification and graph classification by repeating the sampling process 100 times. Each task is executed with five different random seeds, leading to a total of 500 results per task type. We report the mean and standard deviation of these 500 outcomes.
@@
Table 4. Comparison between homophily and non-homophily tasks in pre-training.
@@ 6.2. Performance Evaluation
One-shot performance. We present the results of one-shot node and graph classification tasks on non-homophilic graphs in Tables 1 and 2, respectively. We make the following observations: (1) ProNoG surpasses the vast majority of baseline methods, outperforming the best competitor by up to 21.49% on node classification and 6.50% on graph classification. These results demonstrate its effectiveness in learning prior knowledge from non-homophilic graphs and capturing node-specific patterns. (2) Other graph prompt learning methods, i.e., GPPT, GraphPrompt, and GraphPrompt+, significantly lag behind ProNoG. Their suboptimal performance can be attributed to their inability to account for a variety of node-specific patterns. These results underscore the importance of our conditional prompting in characterizing node embeddings to capture the unique pattern of each node. (3) GPPT is at best comparable to, and often performs worse than other baselines because it is not well suited to few-shot learning.
Few-shot performance. To assess the performance of ProNoG with different amounts of labeled data, we vary the number of shots in the node classification tasks. We present the results in Fig. 3 with several competitive baselines for selected datasets. Note that given the limited number of nodes in Wisconsin, we only conduct tasks up to $3$ shots. We observe that ProNoG generally outperforms these baselines in low-shot scenarios ($k\leq 5$) by a significant margin, showcasing the effectiveness of our approach with limited labeled data. Furthermore, as the number of shots increases, while all methods generally show improved performance, ProNoG remains competitive and often surpasses the other methods. We focus on the one-shot setting for the remaining experimental results.
@@
Figure 4. Results across nodes with varying homophily ratios.
@@ 6.3. Ablation Study
As shown in Table 3, ProNoG consistently outperforms these variants in all but one instance, in which its performance is still competitive. This highlights the necessity of reading out subgraphs with similarity weighting in order to capture the characteristics of each node, and the advantage of using conditional prompt learning to adapt to each node.
@@ 6.4. Analysis on Pre-Training Methods
To investigate the effect of homophily and non-homophily tasks on pre-training, we employ two forms of link prediction from GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) as the homophily tasks, as well as GraphCL (You et al., 2020) and GraphACL (Xiao et al., 2023) as non-homophily tasks. To isolate their effects, we apply the same conditional prompt learning from ProNoG for downstream adaptation.
We present the comparison in Table 4. It can be observed that, for graphs with lower homophily ratios (i.e., Wisconsin and Cornell), non-homophily tasks significantly outperform the homophily tasks. Conversely, for graphs with higher homophily ratios (i.e., PROTEINS and ENZYMES), the performance of homophily and non-homophily tasks becomes more comparable. While the homophily tasks with link prediction may have a slight advantage on highly homophilic graphs, non-homophily tasks are competitive across both homophilic and non-homophilic graphs. Hence, non-homophily tasks provide a more robust solution overall.
6.5. Analysis on Diverse Node Patterns
To evaluate the ability of ProNoG in capturing node-specific patterns, we investigate the classification accuracy of different node groups with varying homophily ratios, i.e., $[0.0,0.2)$, $[0.2,0.4)$, $[0.4,0.6)$, $[0.6,0.8)$, $[0.8,1.0]$. In each node group, we compare the performance of ProNoG with several competitive baselines.
As shown in Fig. 4, ProNoG outperforms the baselines across various node groups, regardless of their unique characteristics reflected in different homophily ratios. These consistent improvements in all groups further demonstrate the effectiveness of ProNoG in capturing diverse node patterns and highlight the advantage of our proposed conditional prompt learning.
@@ 7. Conclusions
In this paper, we explored pre-training and prompt learning on non-homophilic graphs. The goals are twofold: learning comprehensive knowledge irrespective of the varying non-homophilic patterns of graphs, and adapting the nodes with diverse distributions of non-homophily patterns to downstream tasks in a fine-grained, node-wise manner. We first revisit graph pre-training on non-homophilic graphs, providing theoretical insights into the choice of pre-training tasks. Then, for downstream adaptation, we proposed a condition-net to generate a series of prompts conditioned on node-specific non-homophilic patterns. Finally, we conducted extensive experiments on ten public datasets, showing that ProNoG significantly outperforms diverse state-of-the-art baselines.